Forty liters of a 60% disinfection solution are to be mixed with a 10% solution to dilute it to a 20% solution. How much of the 10% solution must be used?
Answer by stanbon(75887)

Forty liters of a 60% disinfection solution are to be mixed with a 10% solution to dilute it to a 20% solution.
How much of the 10% solution must be used?
----------------
Equation:
dis + dis = dis
0.60*40 + 0.10x = 0.20(40+x)
-----
60*40 + 10x = 20*40 + 20x
10x = 40*40
x = 160 liters (amt of 10% solution needed)
===============
Cheers,
Stan H.
==========